OpenStudy (anonymous):

a ball is thrown vertically up with an initial velocity of 60m/s. a) how high will it rise? b) how long will it remain in the air? please help me this is my assignment in physics for fourth year high school please help me...

OpenStudy (anonymous):

to find time V=Vo-gt V=the final velocity=0 g=9.8m/s^2 t=60-0/9.8 t=6.12sec the total time is 2x6.12=12.24 to find how high it will rise y=Vot-1/2gt^2 =60x6.12-1/2x9.8x6.12^2 =183.87m ....i hope this will help u